### v2

# Mitigating Subgroup Disparities in Multi-Label Speech Emotion Recognition: A Pseudo-Labeling and Unsupervised Learning Approach

Abstract
While subgroup disparities and performance bias are increasingly studied in computational research, fairness in categorical Speech Emotion Recognition (SER) remains underexplored. Existing methods often rely on explicit demographic labels, which are difficult to obtain due to privacy concerns. To address this limitation, we introduce an Implicit Demography Inference (IDI) module that leverages pseudo-labeling from a pre-trained model and unsupervised learning using k-means clustering to mitigate bias in SER. Our experiments show that pseudo-labeling IDI reduces subgroup disparities, improving fairness metrics by over 33% with less than a 3% decrease in SER accuracy. Also, the unsupervised IDI yields more than a 26% improvement in fairness metrics with a drop of less than 4% in SER performance. Further analyses reveal that the unsupervised IDI consistently mitigates race and age disparities, demonstrating its potential in scenarios where explicit demographic information is unavailable.
keywords:
Subgroup Disparities, Speech Emotion Recognition, Fairness and Bias
1 Introduction
Speech Emotion Recognition (SER) is essential in human–computer interaction, enhancing user experiences across various applications [1]. As SER becomes increasingly integral to applications ranging from virtual assistants to therapeutic tools, ensuring that these systems operate fairly and without bias is more critical than ever [2, 3]. However, conventional debiasing methods predominantly depend on explicit demographic annotations (e.g., gender, age, race), which are not only challenging to collect due to privacy concerns and ethical issues but are also notably scarce in real-world data [4, 5, 6, 7, 8].
Existing gender debiasing methods in SER are scarce, particularly those that do not require gender labels. Additionally, prior approaches have been designed for single-label categorical SER, while recent research highlights the importance of multi-label SER [9, 10], aligning with psychological studies on emotion co-occurrence [11]. Against these backdrops, our research question is: How can we mitigate subgroup disparities in multi-label SER without relying on explicit demographic annotations?
To address this challenge, we introduce the concept of Implicit Demography Inference (IDI). This novel framework aims to simulate demographic supervision by inferring latent group cues directly from the speech data. The IDI module operates through two techniques: first, by leveraging pseudo-labeling to generate proxy demographic labels using a pre-trained model, and second, by employing unsupervised clustering—specifically, applying K-Means to ECAPA-TDNN embeddings [12]—to uncover inherent group structures. Together, these strategies enable the system to approximate demographic distinctions without requiring manual annotations.
Our experiments on the CREMA-D database [13] provide compelling evidence for the efficacy of this approach. We observe that pseudo-labeling yields improvements in key fairness metrics by over 33%, while the unsupervised clustering approach enhances these metrics by approximately 26%, all with only a minor compromise in overall SER accuracy. Furthermore, the SER model debiased using the unsupervised clustering method can also suppress age and race bias, proving the generalizability of our method. These results underscore the potential of the IDI module to serve as a robust and scalable solution for debiasing in SER, especially in environments where explicit demographic data is limited.
Figure 1: Illustration of proposed debias framework.
2 Related Works
2.1 Debias by Supervised Learning
Supervised learning approaches to mitigating gender bias in SER typically rely on explicitly annotated gender labels during training. Previous studies have implemented fairness constraints or modified training strategies to reduce performance disparities between gender subgroups. Chien et al. [7] employed fairness-constrained adversarial training to develop gender-fair representations, considering both speaker and rater gender to enhance the flexibility of emotion predictions. Similarly, Chien et al. [8] addressed biases originating from both speakers and raters by integrating fairness-aware learning techniques that combine adversarial training with multi-objective optimization, thereby producing gender-neutral emotion predictions. More recently, Upadhyay et al. [14] demonstrated that while SER models may exhibit fairness within their source datasets, significant gender biases can emerge in cross-corpus scenarios. To mitigate this issue, they proposed the Combined Fairness Adaptation method, which leverages both adversarial and contrastive learning to align gender representations across domains, enhancing fairness without compromising recognition performance. However, these studies rely on gender labels, which are often difficult to obtain due to privacy concerns and data limitations. Additionally, they frequently simplify test sets by selecting or removing data to focus on a subset of emotions. Furthermore, these approaches typically treat SER as a single-label classification task, disregarding samples without consensus emotion labels in the test set.
In contrast to these prior studies, our approach retains all data and labels in the test set to provide a more accurate assessment of model performance. Moreover, we define SER as a multi-label task to better capture the complexity of emotion perception, as supported by recent SER research [9, 10, 15] and psychological insights [11]. Emotion perception is inherently multifaceted, with speakers frequently conveying multiple emotions simultaneously. For example, an utterance may express both anger and sarcasm. Thus, multi-label classification presents a more suitable paradigm for SER than traditional single-label classification.
2.2 Debias by Unsupervised Learning
Chou et al. [16] present a method to mitigate individual speaker biases by employing unsupervised clustering on utterance embeddings from a pre-trained speaker verification model. This method effectively clusters utterances with similar characteristics, approximating true speaker identities without explicit identification. The study shows that incorporating these cluster identifiers enables the creation of a fairness-aware SER model that operates equitably across individual speakers. However, this work focuses on dimensional SER—such as arousal, valence, and dominance—and uses a small subset of the original emotion dataset, which, as Ferrer et al. [17] suggest, may lead to overestimated performance in testing. Therefore, we evaluate all experiments using the original test set.
3 Methodology
In this work, we propose a novel framework to mitigate subgroup disparities in gender in multi-label SER without relying on explicit demographic annotations. Central to our approach is the Implicit Demography Inference (IDI) Module, which infers group labels from speech data using a combination of pseudo-labeling and unsupervised clustering. These inferred group labels are then utilized to guide a debias training process. An overview of the proposed framework is illustrated in Fig. 1. In the following sections, we detail the architecture of our proposed IDI module and its integration with debias training techniques.
3.1 Implicit Demography Inference (IDI) Module
The IDI Module is designed to automatically infer latent group labels from speech features through one of two methods: pseudo-labeling and unsupervised clustering.
Pseudo-labeling: we leverage a pre-trained gender detection model [18] [^1] https://huggingface.co/audeering/wav2vec2-large-robust-24-ft-age-gender to generate pseudo gender labels for each input utterance. These labels serve as initial proxies, capturing potential demographic cues inherent in the speech signal. The pseudo-labels offer a straightforward means to incorporate demographic information without manual annotation. The model’s reliability is assessed on the CREMA-D dataset (detailed in Sec. 4.2), achieving an accuracy of 94.4%.
Unsupervised Clustering: We extract embeddings using the ECAPA-TDNN model [12] and apply K-Means clustering [19] (details about parameters are in Appendix Table A1) to these embeddings. The cluster assignments derived from this process are treated as group labels, reflecting latent structures within the data that may correspond to demographic differences. We set the cluster sizes to 2, 4, 8, 16, and 32.
3.2 Debias Training with Inferred Group Labels
Once the group labels are inferred by the IDI Module, they are incorporated into the debias training of the SER model. Specifically, the inferred labels partition the training data into subgroups, allowing us to impose fairness constraints that decrease performance disparities across these groups. We explore four debiasing methods that leverage these inferred group labels:
Reweighting (RW) [20]: This approach adjusts the importance of each sample based on the frequency of its subgroup within each emotion class. The overall weighted loss is computed as:
$\mathcal{L}_{RW}=\sum_{c}\sum_{g}\frac{n_{c}}{n_{c,g}}\,\mathcal{L}_{SER}^{(c,g)},$ (1)
where $n_{c,g}$ is the number of samples in class $c$ belonging to subgroup $g$, reflects the inverse frequency of subgroup $g$ in the specific emotion class. $n_{c}$ is the number of samples in class $c$. $\mathcal{L}_{SER}^{(c,g)}$ denotes the SER loss in class $c$ and subgroup $g$.
Downsampling (DS) [20]: This method balances subgroup distributions by equalizing the number of samples per group within each class. For each class, we determine the minimum number of utterances available among all groups and then randomly select that many samples from each group. This ensures that every subgroup is equally represented during training.
Group Distribution Robust Optimization (GDRO) [21]: GDRO seeks to minimize the worst-case loss among all subgroups, ensuring robust performance across groups. The GDRO objective is formulated as:
$\mathcal{L}_{GDRO}=\max_{g}\,\mathcal{L}_{SER}^{(g)},$ (2)
which directly targets the subgroup with the highest loss.
Group Aware DRO (GADRO) [22]: Extending GDRO, GADRO incorporates an additional regularization term $\frac{\lambda_{GD}}{\sqrt{n_{g}}}$ to penalize large discrepancies between subgroup losses, further promoting fairness. The loss function for GADRO is defined as:
$\mathcal{L}_{GADRO}=\max_{g}\,(\mathcal{L}_{SER}^{(g)}+\frac{\lambda_{GD}}{\sqrt{n_{g}}}),$ (3)
where $\lambda_{GD}$ is a hyperparameter and $n_{g}$ is the group size. We set $\lambda_{GD}$ to 4.
4 Experiments Setup
We select the SSL-based SER model with WavLM base plus [^2] https://huggingface.co/s3prl/converted_ckpts/resolve/main/wavlm_base_plus.pt [23] feature extractor and two linear layers as our primary backbone for the following experiments, because it ranks second on the SER leaderboard [10] while maintaining a lower parameter count compared to top-1 models, XLSR-1B. We use the class-balanced cross-entropy loss [24] as base SER loss $\mathcal{L}_{SER}$.
4.1 Baseline
We compare our proposed methods against several baseline models. Random Baseline: A naive model that assigns predictions uniformly at random, establishing a lower bound on SER performance. Empirical Risk Minimization (ERM): The standard SER model minimizes the overall $\mathcal{L}_{SER}$ on the training data without any fairness constraints. ERM serves as our conventional reference model. Existing Unsupervised Debiasing Methods: We implement state-of-the-art debiasing approaches from computer vision that do not require explicit demographic labels, including LfF [25], and DisEnt [26]. PGDRO: A method that leverages probabilistic pseudo-labels for demography inference [27].
4.2 Dataset
We evaluate our method on the widely used CREMA-D database [13], which contains 7,442 utterances (from 91 actors) in in English across six emotions (anger, disgust, fear, happy, neutral, sad). Each utterance is annotated by at least 7 raters and includes demographic metadata (gender, age, race, and ethnicity). For reproducibility, we use the audio-only annotations and adhere to the EMO-SUPERB splits [10].
Most prior works on debiasing in SER have filtered or selectively sampled data for their experiments. Inspired by Ferrer et al. [17], we simulate biased training data while evaluating models on the original test set to reflect real-world conditions in an unseen setting. To achieve this, we employ ad-hoc data pre-processing to manipulate the training data. First, we assign each sample to a single emotion category based on majority voting. Then, we follow the natural demographic trends in the CREMA-D dataset. In this study, we set the gender imbalance ratio to 1:20 (female to male or male to female). Appendix Table A2 and Table A3 provide a summary of this data distribution.
Table 1: Summary of results in SER accuracy (F1 and ACC) and fairness (TPRgap and DPgap) on gender. Super. specifies the level of supervision for debias. indicates debias with ground truth gender labels. represents the pseudo-labeling approach, and refers to bias unsupervised methods. For unsupervised IDI, we use k-means with k=16. Overall averages the four IDI methods. The Gain represents the improvement compared to the ERM method. $\uparrow$ indicates that higher values correspond to better performance.
[…]
ERM 0.651 0.824 0.278 0.103
Random 0.360 0.509 0.060 0.015
LfF [25] 0.655 0.825 0.268 0.099
DisEnt [26] 0.653 0.825 0.261 0.094
PGDRO [27] 0.643 0.817 0.318 0.121
DS [20] 0.598 0.804 0.094 0.033
DS-IDI 0.580 0.811 0.137 0.045
0.561 0.755 0.201 0.100
RW [20] 0.646 0.823 0.171 0.061
RW-IDI 0.647 0.826 0.197 0.071
0.642 0.817 0.237 0.091
GDRO [21] 0.627 0.796 0.199 0.086
GDRO-IDI 0.632 0.807 0.229 0.088
0.639 0.803 0.228 0.104
GADRO [22] 0.621 0.788 0.185 0.084
GADRO-IDI 0.631 0.807 0.229 0.088
0.638 0.804 0.223 0.098
Overall 0.630 0.808 0.162 0.054
0.631 0.816 0.185 0.059
0.624 0.801 0.205 0.076
[…]
-3.02% -0.95% 33.33% 42.39%
-4.22% -2.77% 26.44% 26.54%
4.3 Evaluation Metrics
For assessing SER, we utilize two primary performance metrics: the macro-F1 score (F1) and Hamming accuracy (ACC). The macro-F1 score is computed by evaluating the F1 measure for each emotion category individually and then averaging the results, thereby handling class imbalances. Hamming accuracy, on the other hand, calculates the ratio of correctly predicted labels to the total number of labels, offering a comprehensive view of overall model performance. In accordance with the approach outlined in [10], we binarize the prediction probabilities using a threshold of $1/|\mathcal{Y}|$, where $\mathcal{Y}$ denotes the set of all emotion classes.
In terms of fairness, our evaluation is based on two critical criteria: Equal Opportunity and Demographic Parity. The Equal Opportunity metric is determined by calculating the root mean square (RMS) of the gap in the true positive rate (TPR) across all emotion classes [28, 29]. This measure ensures that, given the actual positive instances, the likelihood of correctly predicting an emotion does not vary with protected attributes (e.g., gender, race). In addition, Demographic Parity requires that the probability of a positive prediction remains consistent across different demographic groups, independent of the true label. We express this disparity mathematically as:
$DP_{gap}=\sqrt{\sum_{\hat{y}\in\mathcal{Y}}\max_{g\in\mathcal{G}}DP(g,\hat{y})^{2}};$ (4)
$DP(g,\hat{y})=\mathbb{E}\Big{[}P(\hat{y}=1\mid G=g)-P(\hat{y}=1)\Big{]},$ (5)
where $G$ represents a group, $\mathcal{G}$ represents the set of all groups, and $\hat{y}$ represents the emotion prediction. A higher value of DPgap implies a greater imbalance.
5 Results and Analyses
Table 1 presents the averaged results on performance and gender bias across five folds. We evaluate the effectiveness of our proposed methods and provide detailed analyses in the following sections.
5.1 Performance and Effects of Pseudo-labeling
Our results show that pseudo-labeling reduces subgroup disparities in multi-label SER. As seen in Table 1, using pseudo-labels from a pre-trained gender detection model improves the TPRgap fairness metric by 33% and DP by 42% relative to the ERM baseline, with less than a 3% drop in overall accuracy. Compared to PGDRO [27], our modified reweighting strategy with pseudo-labels achieves a better balance between fairness and performance.
5.2 Performance and Effects of Unsupervised Learning
The number of clusters plays a critical role in the effectiveness of this method. Using too few clusters may result in overly coarse groupings that fail to capture the underlying nuances in the data, while employing too many clusters can fragment the data and introduce noise, leading to less meaningful subgroup distinctions. We found that a cluster size of 16 provides a balanced granularity, offering significant improvements in fairness metrics (26% in TPRgap and 27% in DPgap) while maintaining a minor drop in overall performance (a 4% decrease in F1 and a 3% decrease in ACC), as shown in Table 1.
When compared to previously proposed unsupervised debiasing methods such as LfF and DisEnt, our approach exhibits a more favorable trade-off between fairness and SER performance. While these earlier methods yield minor performance improvements, they often incur few debias effects in TPRgap and DPgap. In contrast, our method not only enhances fairness metrics significantly but also maintains higher recognition performance.
To further investigate the underlying reasons behind the method’s effectiveness, we visualize the ECAPA-TDNN embeddings using t-SNE [30], as shown in Fig. 4. The male and female samples are clearly separated, confirming that our clustering captures gender information and supports the scalability and label-free nature of our unsupervised approach.
Table 2: The results of unsupervised clustering on Race and Age. G denotes the number of subgroups within the demographic information. K refers to the number of clusters. Gain represents the performance improvement compared to the baseline model, ERM.
Race (G=3) Age (G=3)
Model K F1 ACC TPRgap DPgap TPRgap DPgap
ERM 0.651 0.824 0.183 0.111 0.153 0.090
DS 16 0.561 0.755 0.142 0.119 0.134 0.095
32 0.575 0.773 0.165 0.112 0.137 0.106
RW 16 0.642 0.817 0.147 0.100 0.140 0.079
32 0.628 0.810 0.172 0.112 0.150 0.095
GADRO 16 0.639 0.803 0.159 0.118 0.138 0.095
32 0.640 0.806 0.162 0.109 0.137 0.096
GDRO 16 0.638 0.804 0.162 0.113 0.122 0.089
32 0.629 0.800 0.172 0.111 0.126 0.090
[…]
32 -5.07% -3.25% 8.33% 0.00% 10.13% -7.50%
Figure 2: Subfigure at left presents the TPRgap while subfigure at right shows the Hamming Accuracy. Bars show the performance of the proposed unsupervised debiasing method evaluated across four techniques under various numbers of clusters.
Figure 3: Illustration of the averaged results across four models trained with the proposed unsupervised method, using different cluster sizes, in terms of Accuracy and TPR.
Figure 4: Illustration of samples in Fold 5 using t-SNE (details about parameters are in Appendix Table A1), with ECAPA-TDNN embeddings and a cluster size of 16. Circles and triangles represent male and female speakers, respectively, while different colors indicate different clusters.
5.3 Potential of Unsupervised Clustering
We assess the applicability of unsupervised clustering to additional demographic attributes, as it does not rely on explicit demographic information. The CREMA-D dataset provides demographic metadata including speakers’ age and race (specifically, Caucasian, African American, and Asian), and their distributions are detailed in Appendix Table A3. In accordance with [31], ages are classified into three subgroups:“Young” (20–35 years),“Middle” (36–49 years), and “Elderly” (50–75 years).
Table 2 presents overall performance across subgroups, demonstrating that the unsupervised method with K=16 clusters improves TPRgap metric by over 12.75%, while maintaining an SER performance decrease of less than 4.76% and 3.55% in F1 and ACC, respectively. This suggests that the method’s performance is robust across different demographic attributes, highlighting its potential for broad applicability for subgroup disparities in multi-label SER tasks.
6 Limitations
While our approach demonstrates robust improvements in mitigating subgroup disparities, several limitations warrant discussion. First, our experiments are based on the CREMA-D dataset, which comprises acted emotional expressions; thus, the generalizability to naturalistic settings remains to be validated. Second, our methods depend heavily on the quality of pre-trained models—the gender detection model for pseudo-labeling and the ECAPA-TDNN for embedding extraction. More pre-trained models can be explored.
7 Conclusion and Future Work
This work empirically examines various debiasing methods in supervised, pseudo-label, and unsupervised learning to mitigate subgroup disparities without explicit demographic information. We leverage pseudo-labels generated by a pre-trained model and introduce an unsupervised clustering approach that integrates with existing debiasing techniques originally designed for demographic labels. Our findings demonstrate that pseudo-label learning improves TPR and DP metrics by 33.33% and 47.90%, respectively, with only a minor 3.02% degradation in SER accuracy. Similarly, our proposed unsupervised method achieves 26.44% and 26.54% improvements in TPR and DP metrics, respectively, with just a 4.22% drop in SER performance. For future work, we will explore more possible bias attributes like content [32], language [33], or speaking style [34].
Supplementary Material
The Supplementary Material provides further details, including Table A1, which outlines the functions for K Means clustering and t-SNE plotting. Additionally, Table A2 presents the data distribution at a 1:20 ratio for Fold 1 of the CREMA-D emotion database, and Table A3 details the distribution of emotions, gender, race, and age across all five folds of the database.
Appendix A Implementation Details
We use the Adam optimizer with a 0.0001 learning rate. The batch size and epoch are set as 32 and 50, respectively. We choose the best models according to the lowest value of the class-balanced cross-entropy loss on the development set. We use two Nvidia Tesla v100 GPUs with 32 GB memory. The total number of GPU hours is around 500.
Table A1: Scikit-learn version is 1.2.2
K Means t-SNE
Parameter Value Parameter Value
init greedy k-means++ init pca
n_clusters [2,4,8,16,32] method barnes_hut
random_state 42 random_state 42
max_iter 1000 n_components 2
batch_size 32 perplexity 40
compute_labels TRUE metric euclidean
reassignment_ratio 0.01 n_iter deprecated
tol 1.00E-04 n_iter_without_progress 300
algorithm lloyd learning_rate auto
Table A2: The table presents the manipulated biased data distribution at a 1:20 ratio in Fold 1.
Emotion Gender Development Test Train
Angry Female 4 102 12
Male 93 58 240
Disgust Female 1 41 4
Male 34 18 86
Fear Female 27 55 127
Male 1 20 6
Happy Female 24 35 102
Male 1 15 5
Neutral Female 19 263 54
Male 384 250 1080
Sad Female 41 35 58
Male 2 7 2
Table A3: Table summarizes detailed data distribution at 1:20 across emotions, gender, and age. AA. means African American. Note that an utterance may be associated with multiple emotions based on the distributional labels, using the threshold, $1/6$. Dev. means Development.
Fold Split Sets Emotion Gender Race Age
Angry Sad Disgust Fear Neutral Happy Male Female Caucasian AA. Asian Young Middle Elderly
1 Train 252 60 90 133 1134 107 1419 357 1070 466 229 927 449 400
Dev. 97 43 35 28 403 25 515 116 476 109 46 365 151 115
Test 160 42 59 75 513 50 574 889 1217 246 0 895 322 246
2 Train 227 77 77 135 1025 111 1282 370 1089 398 165 1028 266 358
Dev. 85 19 31 55 370 32 467 125 349 170 62 188 268 136
Test 147 55 55 69 590 38 896 574 1066 328 76 732 410 328
3 Train 225 105 85 124 1013 113 1276 389 1088 377 200 1059 296 310
Dev. 99 15 28 38 415 23 520 98 470 135 13 334 121 163
Test 143 36 43 96 605 48 820 656 984 328 82 656 574 246
4 Train 243 99 86 141 1036 94 1315 384 1187 390 107 839 511 349
Dev. 67 25 30 38 347 51 428 130 245 159 154 404 56 98
Test 154 30 48 76 609 37 900 574 900 492 82 738 246 490
5 Train 282 78 95 122 1189 80 1504 342 1290 425 117 888 543 415
Dev. 60 36 18 57 262 36 330 139 364 105 0 277 88 104
Test 136 30 48 59 663 60 738 817 817 410 328 1145 246 164